In a test of 30 questions, 20 points are given for each correct answer and 5 points are deducted for each one answered wrong. If Maria answered all of the questions and got a score of 325, how many did she get right?

Answer by nerdybill(6958)

In a test of 30 questions, 20 points are given for each correct answer and 5 points are deducted for each one answered wrong. If Maria answered all of the questions and got a score of 325, how many did she get right?
.
Let x = number she got right
then
30-x = number she got wrong
.
20x - 5(30-x) = 325
20x - 150+5x = 325
25x - 150 = 325
25x = 475
x = 19